OpenStudy (anonymous):

for the function 14/x+3, find (g^-1og)(4). I came up with the inverse y=14-3x/x. (not sure if that is right) and then I thought I had to sub. 4 in for x but that doesn't work. Choices for answers are : 6, 10, 4 or 0.

OpenStudy (anonymous):

Is this the given function?\[g(x)=\frac{ 14 }{ x+3 }\]

OpenStudy (anonymous):

yes, it is a fraction

OpenStudy (anonymous):

If so\[g^{-1}(x)=\frac{ 14-3x }{ x }\]Thus \[(g^{-1}og)(x)=g^{-1}(g(x))=\frac{ 14-3(g(x)) }{ g(x) }\] \[(g^{-1}og)(x)=g^{-1}(g(x))=\frac{ 14-3(g(x)) }{ g(x) }\] \[(g^{-1}og)(4)=g^{-1}(g(4))=\frac{ 14-3(g(4)) }{ g(4) }\] which is 4

OpenStudy (anonymous):

I don't quite see what is happening. What is the inverserse? Once you found the inverse did you sub. 4 for x or is the answer 4 becs. that was what was given to us in question?

OpenStudy (anonymous):

is this what you are saying? 4=14-3(4)=4?

OpenStudy (anonymous):

There is a postulate if you want know either, it says \[(f^{-1}of)(x)=x\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did I get the inverse correct

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok, after finding the inverse do I sub 4 in for x?

OpenStudy (anonymous):

No sub g(x) in the place of x of the inverse function

OpenStudy (anonymous):

ok what is g(x)? what is done with the 4 that is given

OpenStudy (anonymous):

is g(x) the orginial equation?

OpenStudy (anonymous):

I mean g(4)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so I need to solve for this equation to get my answer ? y={14-3(14/x+3)}/14/x+3 what happens to the 4 in the given question?

OpenStudy (anonymous):

yes, then insert 4 in place x here.

OpenStudy (anonymous):

oh--okay I get it. Thanks

OpenStudy (anonymous):

so what was thee Answer? please...

OpenStudy (anonymous):

so the right answer is 4